OpenStudy (anonymous):

There are (3^5)2X3^0 leaves on a tree. What is the total number of leaves on the tree? 3^3 3^7 3^10 3^25

OpenStudy (aum):

(3^5)2X3^0 is not clear. Can you retype it clearly or post a screenshot of the question?

OpenStudy (gorv):

3^5)^2=3^(5*2)=3^10

OpenStudy (gorv):

3^0=1

OpenStudy (anonymous):

thx almost second guessed myself

OpenStudy (triciaal):

(3^5)2X3^0 this is not clear